A charge Qa = 4 X 10^-9 C experiences a et electrostatic force of 1.2 X10^-6 N toward the north.

a. Find the electric field (magnitude and direction) at the location of Qa.

b. If the charge Qa is replaced by a charge Qb = -1.5 X10^-9 C, what force (magnitude and direction) will Qb experience?
Answer #1

a) Electric Field E= F/Q = 1.2*10-6/(4*10-9) =300 N/c along north
b) Electric Force experienced by Qb = (1.2*10-6)*(1.5*10-9)/(4*10-9) =4.5*10-7 N towards South
  Electric Field E= F/Q = 4.5*10-6/(1.5*10-9) =300N/c along north